OpenStudy (anonymous):

If a line with a slope of -2 crosses the y-axis at (0, 3), what is the equation of the line? y − 2x = 3 -2y + 3x = 1 y + 3x = 2 y + 2x = 3

OpenStudy (whpalmer4):

The point-slope formula allows you to write the equation of a line with slope \(m\) passing through a point \((x_0,y_0)\): \[y-y_0 = m(x-x_0)\] You have \(m = -2\) and \((x_0,y_0) = (0,3)\) You will have to do a little bit of algebra after writing the equation to get \(x\) and \(y\) on the same side and the numbers combined on the other side.